OpenStudy (anonymous):

find a formula for the inverse of the function: F(x)=e^x^3 I'm extremely confused on the this problem...

OpenStudy (anonymous):

To find an inverse, all you need to do is switch the x's and y's, and then solve for y. So here, we're going to have \[\begin{align} y=e^{x^3}\\ x=e^{y^3}\\ \ln x = y^3\\ y=\sqrt[3]{\ln x} \end{align}\] will be your inverse function.

OpenStudy (anonymous):

Oh...I get it now. thanks so much!